# Generalization performance of narrow one-hidden layer networks in the teacher-student setting

### v3

@@ 3 Numerical experiments and comparison with theoretical predictions
Figure 2: Generalization error $\epsilon_{g}$ and teacher-student overlap for a regression task with MSE loss without noise ($\Delta^{\star}=0$), as functions of the sample rate $\tilde{\alpha}$. The inverse temperature is $\beta=10$, weight decay $\lambda=0.1$, and the activations functions are (Left) $\sigma(x)={\rm ReLU}(x)$, (Right) $\sigma(x)={\rm Erf}(x/\sqrt{2})$. Theoretical curves are obtained from the RS equations. The diagonal and off-diagonal teacher-student overlaps are denoted by $m_{0}^{*}=m_{d}^{*}+m_{a}^{*}/K$ and $m_{1}^{*}=m_{a}^{*}/K$, respectively, with their values shown on the right axis. The solid line represents the equilibrium solution of the RS equations, i.e. the one which globally maximizes the RS free entropy. The specialization transition beyond which the specialized solution is the equilibrium is denoted $\tilde{\alpha}_{\rm spec}$. For $\mathrm{Erf}$, the dashed line represent the metastable solution of the RS equations different from the equilibrium. There, a second transition appears at $\tilde{\alpha}_{s}<\tilde{\alpha}_{\rm spec}$, where the specialization branch starts existing. Each dot averages 500 LD samples for $K=10$, $N=500$ for a single dataset. Generalization error is computed on 1000 test samples; error bars show standard deviations.
@@
As our analysis describes a Bayesian posterior over the first-layer weights with a $\beta$-dependent likelihood, our primary algorithm for testing is Langevin dynamics (LD), due to its explicit temperature dependence and theoretical guarantee of sampling the Gibbs measure when it reaches equilibrium. We consider two distinct initializations: LD Planted Init, where the student parameters are initialized near the ground truth, and LD Random Init, with i.i.d. $\mathcal{N}(0,1)$ initial weights. Across data regimes, at least one of these LD variants consistently displays a test error matching well the theoretical prediction extracted from a solution of the RS equations that locally maximizes $\Phi_{\beta,K}^{\rm RS}$. The global maximizer corresponds to the “equilibrium branch” of the theory, describing typical student weights dominating the Gibbs measure, which may coexist with other solutions associated with metastable states, as described below. We also compare our $\beta\to\infty$ theory to gradient descent (GD), identifying regimes where GD performance aligns well with the theory, despite lacking guarantees for solving the non-convex ERM problem. Although neither GD nor LD consistently reach the minimal generalization error of ERM across all $\tilde{\alpha}$, our theoretical framework reliably describes their performance. A single GD run took at most $16$ h, while LD required up to $58$ h, including GPU and CPU time. Total compute time for all figures is $42$ effective GPU hours (see SM).
@@
While our theory extends to broader settings, for the sake of illustration we focus on regression with a mean-square (MSE) loss and classification with Hinge loss. Given the non-convex nature of the problem, our theory yields multiple stationary solutions of the RS free entropy potential, characterized by the order parameters $(m^{*},q^{*},v^{*})$. We describe them in the next paragraph and Fig. 1. Once this picture is further detailed, we then provide numerical validations for all predicted observables for regression at finite $\beta$ (Fig. 2) and ERM (Fig. 3). Finally, we numerically validate our theory adapted to study the Bayes-optimal (BO) setting for classification (Fig. 4), and later use it to compare BO with ERM for regression across different weight decay values $\lambda$ (Fig. 5).
@@
Figure 3: Generalization error $\epsilon_{g}$ for regression with MSE loss and no noise ($\Delta^{\star}=0$), as functions of the sample complexity $\tilde{\alpha}$. (Left) $\sigma(x)=\mathrm{ReLU}(x)$; (Right) $\sigma(x)=\mathrm{Erf}(x/\sqrt{2})$. Results corresponds to ERM ($\beta\to\infty$) with weight decay $\lambda=0.01$, where typical configurations are sampled from the memorization state when $\tilde{\alpha}<\tilde{\alpha}_{\rm spec}$. The equilibrium branch is the solid black line, the metastable one is dashed and the orange is the Bayes-optimal error, which vanishes for $\tilde{\alpha}\geq 1$. Numerical results are for a single dataset with $K=10$ and $N=500$, using LD at $\beta=10^{4}$ (averaged over $500$ equilibrium samples) and a single GD estimator obtained from one initialization. Generalization error is computed on 1000 test samples; error bars indicate standard deviation.
In Fig. 2, we illustrate the remarkable agreement between the analytical results for the overlaps, Eq. (13), and the generalization error described in Sec. 2.2, with numerical results obtained from LD. Notably, systems with as few as $K=10$ hidden units are already well described by our large-width theory.
@@
Previous work using a simplified annealed calculation [oostwal_hidden_2021] showed that the choice of activation function influences the nature of the learning phase transition. In a similar regression setting, the authors found that networks with ReLU activation undergo a continuous phase transition at a critical value of $\tilde{\alpha}$. We recover this result, as shown in the left panel in Fig. 2, where we observe a specialization transition at $\tilde{\alpha}_{\rm spec}$. However, our more accurate quenched computation predicts that for sufficiently large $\beta$, i.e. close to the ERM setting, ReLU networks exhibit a discontinuous phase transition (see left panel in Fig. 3). In contrast, for networks with Erf activation we observe a discontinuous transition independently of $\beta$ (right panel on Fig. 2). Discontinuous transitions are associated with the presence of metastable thermodynamic states. Depending on the initialization, these states can trap LD, preventing it from sampling the Gibbs measure. We identify the presence of metastability by running LD with both random and teacher-close initializations of the weights. In the latter case, LD follows the specialized branch for $\tilde{\alpha}\geq\tilde{\alpha}_{s}$, where the teacher-student overlap increases with the training set size. Conversely, with random initialization, LD approaches the specialization branch at a value $\tilde{\alpha}>\tilde{\alpha}_{\rm spec}$ and not at the theoretical transition. This signals the possible presence of a hard phase for LD, where it fails to equilibrate when the weights are initialized randomly. A theoretical explanation for this may require a finer analysis taking into account the effects of “replica symmetry breaking” in physics jargon [mezard1987spin, antenucci2019glassy].
Which branch becomes stable at the specialization transition $\tilde{\alpha}_{\rm spec}$ depends on $\beta$. At intermediate temperatures (e.g. $\beta=10$), the transition is between the permutation-symmetric and the specialized branches (Fig. 2). In contrast, in the $\beta\to\infty$ limit, the transition occurs between the memorization and specialized branches, indicating that at small temperatures the PS branch loses stability to the memorization branch. The results obtained using LD with random initializations support this, as the generalization error and the order parameters converge to the corresponding memorization branch for $\tilde{\alpha}<\tilde{\alpha}_{\rm spec}$ in the $\beta\to\infty$ limit (Fig. 3). Notably, we observe in the same figure that GD with random initialization recovers the statistics of typical ERM minimizers. Such observation is consistent with recent work showing that (S)GD concentrates on solutions with probabilities close to the Bayesian posterior [hennick2025almost, mingard2021sgd, smith2018stochastic, JMLR:v18:17-214], providing a possible explanation for why GD reliably finds typical ERM minimizers despite nonconvexity. A similiar behavior was observed for two-layer neural networks in the extensive width regime in the Bayes-optimal setting with $\sigma(x)=x^{2}$ [maillard_quadratic]  [^1] Maillard et al. [maillard_quadratic] observed a gap between the performance of GD with small regularization and the Bayes-optimal one in the noisy teacher-student regression setting with $\sigma(x)=x^{2}$. Interestingly, in their paper GD follows a smooth trajectory, which may reflect the memorization branch we described here, a branch that could still exist in the extensive-width regime $K=O(N)$ they consider.. However, we observed that GD reaches the specialization branch at a larger value than $\tilde{\alpha}_{\rm spec}$: specifically, at $\tilde{\alpha}^{\text{ReLU}}_{\rm GD}\simeq 2.3$ and $\tilde{\alpha}^{\text{Erf}}_{\rm GD}\simeq 2.5$, respectively. This difference may be related to the existence of other subdominant, atypical configurations attracting GD not described by our theory. As it should, we observe that the Bayes-optimal error (BOE) [aubin_committee_2018] is a lower bound for ERM, which can however approach it closely when tuning the regularization $\lambda$ properly, see the last paragraph of this section.
@@
Analogous comparisons between theory and simulations can be done in the classification case (where $f=\mathrm{sign}$). In Fig. 4 we show the accuracy ($1-\epsilon_{g}$) as a function of $\tilde{\alpha}$ for both ReLU and Erf activations in the noiseless case ($\Delta^{\star}=0$) found by doing ERM on the hinge loss, in the BO setting (obtained by setting $\lambda\beta=1$ and $\beta\to\infty$). In both cases the theory predicts that the equilibrium is given by the specialized branch. This is in contrast with the $\mathrm{sign}$ activation, which was investigated in [schwarze_learning_1993, schwarze_discontinuous_1993], for which one has a discontinuous transition from the permutation-symmetric to the specialized branch. A study of the small data regime where the size of the dataset is proportional to the size of the input (i.e. $\alpha=P/N=O(1)$), shows that in both those cases the transition from the PS solution is continuous (see SM for additional plots and details).
@@
The optimal regularization strength depends on the sample complexity $\tilde{\alpha}$, reducing overfitting in low-data regimes while enabling alignment with the teacher in high-data regimes. Contrary to convex optimization problems [krogh1991simple], the replica analysis reveals a discontinuous behaviour of the optimal regularization as a function of the training set size, a behaviour generally difficult to grasp using statistical learning bounds. In particular, for $\tilde{\alpha}<\tilde{\alpha}^{\mathrm{BO}}_{\mathrm{sp}}$ there coexist two learning phases: Memorization and PS. Although neither describes a phase where the student aligns with the teacher, the latter is still preferable in terms of generalization performance since it matches the BO performance for small $\tilde{\alpha}$. In this regime, a large regularization strength is optimal: it keeps the system in the PS phase, avoiding overfitting and preventing harmful memorization.
@@
Figure 5: Phase diagram (Left) and generalization error (Right) of a two-layer student neural network with activation function $\mathrm{Erf}(x/\sqrt{2})$, trained on a regression task via ERM ($\beta\to\infty$) with MSE loss $\ell(x,y)=(x-y)^{2}/2$, as a function of the inverse regularization $\lambda^{-1}$ and the sample complexity $\tilde{\alpha}$. The training data are generated by a teacher network with matching architecture, label noise $\Delta^{*}=10^{-4}$, and $K=10$ hidden units. (Left) Phase diagram in the $(\tilde{\alpha},\lambda^{-1})$ plane showing three learning regimes. In the permutation-symmetric (PS) phase, hidden units across different student networks—initialized differently but trained on data generated by the same teacher—remain uncorrelated ($q_{d}^{*}\simeq 0$); this occurs for small $\tilde{\alpha}$ and $\lambda>\lambda_{m}$. In the memorization phase, hidden units across different student networks—initialized differently but trained on data generated by the same teacher—become mutually correlated, yet remain uncorrelated with the teacher’s hidden units ($q_{d}^{*}>0$, $m_{d}^{*}\simeq 0$); this corresponds to intermediate $\tilde{\alpha}$ and $\lambda<\lambda_{m}$. The bold blue line at $\tilde{\alpha}_{m}$ marks the onset of stability of this solution. In the specialization phase, student units align with teacher units ($m_{d}^{*}>0$). The bold green line $\tilde{\alpha}_{sp}$ marks the stability threshold, while the dashed purple line $\tilde{\alpha}_{s}$ indicates where the specialized solution first appears (metastable until crossing the green boundary). The dashed vertical red line $\tilde{\alpha}^{BO}_{sp}$ is the Bayes-optimal specialization threshold, and the horizontal dashed pink line $\lambda_{m}$ marks where the regularization becomes beneficial as a network can escape to go from the PS to memorization region. Bold brown indicates that these solutions are present only as metastable states. (Right) Bayesian optimal error (BOE, orange) and Gibbs generalization error (red and purple tones) as functions of $\tilde{\alpha}$. Bold curves denote equilibrium solutions, while dashed curves represent metastable ones. Across different $\tilde{\alpha}$ (highlighted in red in both panels), $\lambda_{\mathrm{opt}}$ denotes the value of the regularization strength for which the student’s generalization error best approaches the BOE.

### v4

@@ 3 Numerical experiments and comparison with theoretical predictions
As our analysis describes a Bayesian posterior over the first-layer weights with a $\beta$-dependent likelihood, our primary algorithm for testing is Langevin dynamics (LD), due to its explicit temperature dependence and theoretical guarantee of sampling the Gibbs measure when it reaches equilibrium. We consider two distinct initializations: LD Planted Init, where the student parameters are initialized near the ground truth, and LD Random Init, with i.i.d. $\mathcal{N}(0,1)$ initial weights. Across data regimes, at least one of these LD variants consistently displays a test error matching well the theoretical prediction extracted from a solution of the RS equations that locally maximizes $\Phi_{\beta,K}^{\rm RS}$.
Figure 2: Generalization error $\epsilon_{g}$ and teacher-student overlap for noise-free regression task ($\Delta^{\star}=0$) as functions of the sample-to-parameter ration $\tilde{\alpha}$. Results are shown for inverse temperature $\beta=10$, weight decay $\lambda=0.1$, and activation functions $\sigma(x)={\rm ReLU}(x)$ (Left), which exhibit a continuous specialization transition, and $\sigma(x)={\rm Erf}(x/\sqrt{2})$ (Right) which exhibit a discontinuous specialization transition. Theoretical curves are derived from the RS equations. The teacher-student overlaps are parametrized as $m_{0}^{*}=m_{d}^{*}+m_{a}^{*}/K$ (diagonal) and $m_{1}^{*}=m_{a}^{*}/K$ (off-diagonal). The solid lines represent the equilibrium RS solutions (globally maximizing free entropy), while the dashed lines for the Erf case denote metastable solutions. The specialized branch becomes the equilibrium at $\tilde{\alpha}\geq\tilde{\alpha}_{\rm spec}$; for Erf, this branch emerges at a spinodal point $\tilde{\alpha}_{s}<\tilde{\alpha}_{\rm spec}$. Markers represent averages over 500 LD samples with $K=10$ and $N=500$, for a single dataset. Generalization error is computed on 1000 test samples; error bars show standard deviations. Notably, for $\tilde{\alpha}>\tilde{\alpha}_{\rm spec}$, LD with random initialization fails to follow the RS equilibrium branch, suggesting metastability effects or dynamical barriers not fully captured by the RS equilibrium analysis.
We perform numerical simulations at $K=10$ and $N=500$. While the theoretical equations predict a slight shift in the specialization threshold as $K$ increases towards the asymptotic limit, $K=10$ and $N=500$ is sufficient to capture the relevant physical features—such as the specialization transition and the emergence of a hard phase—while allowing for extensive LD sampling within reasonable computational limits.
The global maximizer corresponds to the “equilibrium branch” of the theory, describing typical student weights dominating the Gibbs measure, which may coexist with other solutions associated with metastable states, as described below. We also compare our $\beta\to\infty$ theory to gradient descent (GD), identifying regimes where GD performance aligns well with the theory, despite lacking guarantees for solving the non-convex ERM problem. Although neither GD nor LD consistently reach the minimal generalization error of ERM across all $\tilde{\alpha}$, our theoretical framework reliably describes their performance. A single GD run took at most $16$ h, while LD required up to $58$ h, including GPU and CPU time. Total compute time for all figures is $42$ effective GPU hours (see SM).
@@
While our theory extends to broader settings, for the sake of illustration we focus on regression with a mean-square (MSE) loss and classification with Hinge loss. Given the non-convex nature of the problem, our theory yields multiple stationary solutions of the RS free entropy potential, characterized by the order parameters $(m^{*},q^{*},v^{*})$. We describe them in the next paragraph and Fig. 1. Once this picture is further detailed, we then provide numerical validations for all predicted observables for regression at finite $\beta$ (Fig. 2) and ERM (Fig. 3). Finally, we numerically validate our theory adapted to study the Bayes-optimal (BO) setting for classification (Fig. 4).
@@
Figure 3: Generalization error $\epsilon_{g}$ for noise-free regression ($\Delta^{\star}=0$) as a function of the sample-to-parameter ratio $\tilde{\alpha}$. (Left) $\sigma(x)=\mathrm{ReLU}(x)$; (Right) $\sigma(x)=\mathrm{Erf}(x/\sqrt{2})$. At large $\beta$ both activations exhibit a discontinuous specialization transition. Results corresponds to ERM ($\beta\to\infty$ limit) with weight decay $\lambda=0.01$, where typical configurations are sampled from the memorization state for $\tilde{\alpha}<\tilde{\alpha}_{\rm spec}$. The equilibrium and metastable RS branches are denoted by solid and dashed black lines, respectively, while the orange line represents the Bayes-optimal error (BOE), which vanishes for $\tilde{\alpha}\geq 1$. Numerical results for a single dataset with $K=10$ and $N=500$, are obtained via LD at $\beta=10^{4}$ (averaging over $500$ samples) and a single GD estimator with random initialization. Generalization error is computed on 1000 test samples; error bars indicate standard deviation. Notably, for $\tilde{\alpha}<\tilde{\alpha}_{\rm spec}$, LD results whith planted initialized can fall below the Bayes-optimal prediction. This apparent violation is likely due to initialization bias and slow relaxation dynamics, which prevent the system from reaching the true equilibrium within the simulation time.
@@
In Fig. 2, we illustrate the remarkable agreement between the analytical results for the overlaps, Eq. (13), and the generalization error described in Sec. 2.2, with numerical results obtained from LD. Notably, systems with as few as $K=10$ hidden units are already well described by our large-width theory. Previous work using a simplified annealed calculation [35] showed that the choice of activation function influences the nature of the learning phase transition. In a similar regression setting, the authors found that networks with ReLU activation undergo a continuous phase transition at a critical value of $\tilde{\alpha}$. We recover this result, as shown in the upper-left panel in Fig. 2, where we observe a specialization transition at $\tilde{\alpha}_{\rm spec}$. In this large-dataset regime, the quenched computation effectively reproduces the annealed prediction, which corresponds to the small $\beta$ limit. However, our more accurate quenched computation predicts that for sufficiently large $\beta$, i.e. close to the ERM setting, ReLU networks exhibit a discontinuous phase transition (see left panel in Fig. 3). In contrast, for networks with an Erf activation, we observe a discontinuous transition, independently of $\beta$ (upper-right panel of Fig. 2 and right panel in Fig. 3). Discontinuous transitions are associated with the presence of metastable thermodynamic states. Depending on the initialization, these states can trap LD, preventing it from sampling the Gibbs measure. We identify the presence of metastability by running LD with both random and teacher-close initializations of the weights. In the latter case, LD follows the specialized branch for $\tilde{\alpha}\geq\tilde{\alpha}_{s}$, where the teacher-student overlap increases with the training set size. Conversely, with random initialization, LD approaches the specialization branch at a value $\tilde{\alpha}>\tilde{\alpha}_{\rm spec}$ and not at the theoretical transition. This signals the possible presence of a hard phase for LD, where it fails to equilibrate when the weights are initialized randomly. A theoretical explanation for this may require a finer analysis taking into account the effects of “replica symmetry breaking” in physics jargon [27, 5].
The specific branches that becomes stable around the specialization transition $\tilde{\alpha}_{\rm spec}$ depends on $\beta$. At intermediate temperatures (e.g. $\beta=10$), the transition occurs between the permutation-symmetric and specialized branches (see right-upper panel of Fig. 2). In contrast, in the $\beta\to\infty$ limit, the transition shifts to one between the memorization and specialized branches (see right panel of Fig. 3). This indicates that at low temperatures the PS branch loses stability to the memorization branch.
The results obtained using LD with random initializations support this; in the $\beta\to\infty$ limit, both the generalization error and order parameters converge to the corresponding memorization branch for $\tilde{\alpha}<\tilde{\alpha}_{\rm spec}$ (Fig. 3). Notably, we observe in the same figure that GD with random initialization recovers the statistics of typical ERM minimizers. Such observation aligns with recent work suggesting that (S)GD concentrates on solutions with probabilities close to the Bayesian posterior [21, 28, 47, 26], providing a possible explanation for why GD reliably finds typical ERM minimizers despite non-convex landscape. However, at finite system size, initializing at the planted solution alongside incomplete equilibration may trap the dynamics in atypical low-error states. Consequently, the observed generalization error can dip below the asymptotic Bayes-optimal bound, reflecting the lack of full posterior sampling.
A similar behavior was observed for two-layer neural networks in the extensive-width regime in the Bayes-optimal setting with $\sigma(x)=x^{2}$ [25]  [^1] Maillard et al. [25] observed a gap between the performance of GD with small regularization and the Bayes-optimal one in the noisy teacher-student regression setting with $\sigma(x)=x^{2}$. Interestingly, in their paper GD follows a smooth trajectory, which may reflect the memorization branch we described here, a branch that could still exist in the extensive-width regime $K=O(N)$ they consider.. In our case, we observed that GD reaches the specialization branch at values significantly larger than $\tilde{\alpha}_{\rm spec}$– specifically, at $\tilde{\alpha}^{\text{ReLU}}_{\rm GD}\simeq 2.3$ and $\tilde{\alpha}^{\text{Erf}}_{\rm GD}\simeq 2.5$. This discrepancy likely stems from subdominant, atypical configurations that attract GD but are not captured by our theory. While the BOE [6] remains a strict lower bound in the thermodynamic limit, finite-size effects and biased (planted) initialization can lead to apparent violations in LD simulations at small $\tilde{\alpha}$ due to incomplete equilibration. Nevertheless, as discussed in the next section, ERM can closely approach the BOE when the regularization $\lambda$ is properly tuned.
@@
Analogous comparisons between theory and simulations can be done in the classification case (where $f=\mathrm{sign}$). In Fig. 4 we show the accuracy ($1-\epsilon_{g}$) as a function of $\tilde{\alpha}$ for both ReLU and Erf activations in the noiseless case ($\Delta^{\star}=0$) found by doing ERM on the hinge loss, in the BO setting (obtained by setting $\lambda\beta=1$ and $\beta\to\infty$). In both cases the theory predicts that the equilibrium is given by the specialized branch. This is in contrast with the $\mathrm{sign}$ activation, which was investigated in [43, 45], for which one has a discontinuous transition from the permutation-symmetric to the specialized branch. A study of the small data regime where the size of the dataset is proportional to the size of the input (i.e. $\alpha=P/N=O(1)$), shows that in both those cases the transition from the PS solution is continuous (see SM for additional plots and details).
4 Phase diagram and extensive-width limit
Understanding how the regularization strength and the data-to-parameter ratio shape the equilibrium solution is a central question in learning theory. In this section, we characterize the resulting phase diagram and investigate how the limit $K\to\infty$ in our setting relates to the $N\to\infty$ limit, clarifying the interplay between width and input dimension.
Figure 5: Phase diagram (Left) and generalization error (Right) of a two-layer student neural network with activation function $\mathrm{Erf}(x/\sqrt{2})$, trained on a regression task via ERM ($\beta\to\infty$) with MSE loss $\ell(x,y)=(x-y)^{2}/2$, as a function of the inverse regularization $\lambda^{-1}$ and the sample-to-parameter ration $\tilde{\alpha}$. The training data are generated by a teacher network with matching architecture, label noise $\Delta^{*}=10^{-4}$, and $K=10$ hidden units. (Left) Phase diagram in the $(\tilde{\alpha},\lambda^{-1})$ plane showing three learning regimes. In the permutation-symmetric (PS) phase, hidden units across different student networks—initialized differently but trained on data generated by the same teacher—remain uncorrelated ($q_{d}^{*}\simeq 0$); this occurs for small $\tilde{\alpha}$ and $\lambda>\lambda_{m}$. In the memorization phase, hidden units across different student networks—initialized differently but trained on data generated by the same teacher—become mutually correlated, yet remain uncorrelated with the teacher’s hidden units ($q_{d}^{*}>0$, $m_{d}^{*}\simeq 0$); this corresponds to intermediate $\tilde{\alpha}$ and $\lambda<\lambda_{m}$. The bold blue line at $\tilde{\alpha}_{m}$ marks the onset of stability of this solution. In the specialization phase, student units align with teacher units ($m_{d}^{*}>0$). The bold green line $\tilde{\alpha}_{sp}$ marks the stability threshold, while the dashed purple line $\tilde{\alpha}_{s}$ indicates where the specialized solution first appears (metastable until crossing the green boundary). The dashed vertical red line $\tilde{\alpha}^{BO}_{sp}$ is the Bayes-optimal specialization threshold, and the horizontal dashed pink line $\lambda_{m}$ marks where the regularization becomes beneficial as a network can escape to go from the PS to memorization region. Bold purple scripts indicates that these solutions are present only as metastable states. (Right) Bayesian optimal error (BOE, orange) and Gibbs generalization error as functions of $\tilde{\alpha}$. Solid curves denote equilibrium solutions, while dashed curves represent metastable ones. Across different $\tilde{\alpha}$ (highlighted in red in both panels), $\lambda_{\mathrm{opt}}$ denotes the value of the regularization strength for which the student’s generalization error best approaches the BOE.
@@ 4 Phase diagram and extensive-width limit
The optimal regularization strength depends on the sample-to-parameter ration $\tilde{\alpha}$, reducing overfitting in low-data regimes while enabling alignment with the teacher in high-data regimes. Contrary to convex optimization problems [23], the replica analysis reveals a discontinuous behaviour of the optimal regularization as a function of the training set size, a behaviour generally difficult to grasp using statistical learning bounds. In particular, for $\tilde{\alpha}<\tilde{\alpha}^{\mathrm{BO}}_{\mathrm{sp}}$ there coexist two learning phases: Memorization and PS. Although neither describes a phase where the student aligns with the teacher, the latter is still preferable in terms of generalization performance since it matches the BO performance for small $\tilde{\alpha}$. In this regime, a large regularization strength is optimal: it keeps the system in the PS phase, avoiding overfitting and preventing harmful memorization.
@@
The extensive width limit—
It is natural to ask how the following two limits are related: the one treated in the present paper, namely $K\to\infty$ after $N\to\infty$, and the small-but-extensive width limit considered in [8], where $N,K\to\infty$ jointly with $K=\gamma N$ and $\gamma\ll 1$ but independent of $N$. In Fig. 6, we address this question in the Bayes-optimal setting by plotting the Bayes-optimal generalization error for three activation functions, from left to right:
$\sigma(x)=\mathrm{Erf}\!\left(\frac{x}{\sqrt{2}}\right),\qquad\sigma(x)={\rm He}_{3}(x),\qquad\sigma(x)={\rm He}_{2}(x)+\frac{1}{\sqrt{6}}{\rm He}_{3}(x),$
where we use the normalized probabilist’s Hermite polynomials ${\rm He}_{n}(x):=\frac{(-1)^{n}}{\sqrt{n!}}\,e^{x^{2}/2}\,\frac{d^{n}}{dx^{n}}e^{-x^{2}/2}.$ Across the panels, the solid lines for different values of $\gamma$ correspond to the extensive-width prediction of [8], while the dashed lines represent the results obtained from our theory of narrow-network. In the left and middle panels of Fig. 6, the two theories coincide for all values of the sample-to-parameter ratio $\tilde{\alpha}$. In these cases, the collapse of the predictions for different $\gamma$ in [8] is a consequence of the absence of the second Hermite coefficient in the activation functions. In particular, the specialization transition matches that of the extensive-width case. In the right panel, we observe that as $\gamma\to 0$ the two theories agree across all data-to-parameter ratios (here, taking $\gamma=1/50$ already yields a quasi-perfect match), while for non-vanishing $\gamma$ they (quickly) converge to the same limit as $\tilde{\alpha}\to\infty$. These observations suggest that the large-$N$ and large-$K$ limits commute for any $\tilde{\alpha}$ when the activation function has zero second Hermite coefficient, or “quasi commute” when $\tilde{\alpha}$ diverges.
Figure 6: Comparison of the generalization error $\epsilon_{g}-\Delta^{*}$ as a function of the sample-to-parameter ratio $\tilde{\alpha}=P/(NK)$ in the Bayes-optimal setting, between our theory for narrow-networks (dashed black line, corresponding to the limit $K\to\infty$ taken after $N\to\infty$) and the extensive-width theory of [8] (colored lines, corresponding to the limit $K,N\to\infty$ with $K=\gamma N$). (Left) $\sigma(x)=\mathrm{Erf}(x/\sqrt{2})$ and $\Delta^{*}=0.005$, (Middle) $\sigma(x)={\rm He}_{3}(x)$ and $\Delta^{*}=0.005$, (Right) $\sigma(x)={\rm He}_{2}(x)+\frac{1}{\sqrt{6}}{\rm He}_{3}(x)$ and $\Delta^{*}=0.1$.